# Ultrafast dynamics of moments in bulk ferromagnets

### v1

@@ I Introduction
Tranchida et al.  [47] have formulated a rigorous and systematic statistical averaging procedure, based on functional calculus and incorporating a few microscopic material parameters. This procedure can be applied to derive set of LLB-like differential equations for the first and second moments of spin variables. The differential equation governing the dynamics of the first moment preserves the original form of the LLB equation and inherently generates a thermal longitudinal damping. Damping is calculated on the basis of variations in magnetic moments due to thermal noise, thus avoiding the need to provide prior information on longitudinal and transverse thermal damping. This equation also includes a second moment that does not factorize in a sum of first moments, suggesting that the magnetization variance may also fluctuates over time. In this model called dynamical LLB (dLLB), the damping torque that influences the precession of the magnetization, is governed by the dynamics of both the first and second moments. In return, the second moment also fluctuates in time and is itself the solution of a differential equation. The entire differential system constitutes an open hierarchy that can be closed in various ways [47] and solved numerically in a consistent manner.
@@ II Theoretical framework
To accurately simulate the ultrafast magnetization dynamics caused by laser absorption, an essential component in atomistic simulations of magnetization dynamics is needed [49]. The so-called phenomenological three-temperature model (3TM) [24] considers three interacting reservoirs, namely hot electrons, phonons and spins, and their corresponding temperatures, that exchange heat over time [50]. The electron reservoir is affected firstly by the laser pump, due to the efficiency of the Lorentz force produced on electrons by strong electromagnetic fields. Then, the spins and the phonons become thermally unbalanced, and their temperatures change at a rate defined by their heat capacities and the strength of their coupling to the electron reservoir, as well as interactions among themselves.
@@
$\bm{\omega}_{i}$ $\equiv\sum_{j\neq i}J_{ij}{\bm{s}}_{j}.$ (1)
This allows the formation of a skew-symmetric matrix $\Omega$ where $\Omega_{IJ}\equiv\epsilon_{IJK}\omega_{i}^{K}$ are its components. We also set $\Sigma_{i}\equiv\langle{\bm{S}}_{i}\otimes{\bm{S}}_{i}\rangle$ a symmetric $3\times 3$ matrix, a symmetric matrix $\Gamma_{i}\equiv{\bm{s}}_{i}\otimes\bm{s}_{i}$ and a (general) matrix $M_{i}\equiv{\bm{s}}_{i}\otimes{\bm{\omega}}_{i}$, which can be decomposed into a symmetric (resp. antisymmetric) part labelled ${M}_{i}^{S}$ (resp.${M}_{i}^{A}$). The dLLB model is reduced to the following set of equations
@@
where $\alpha$ is the dimensionless transverse Gilbert damping parameter and ${\sf{Tr}}$ stands for the trace of the matrix. In eq. (3), the commutator of two matrices is appearing as an operator that produces a matrix by contracting its internal index, thereby reducing it to a symmetric matrix, as expected. Equations (2) and (3) are numerically integrated over time by a global Runge-Kutta scheme of 4-th order, with an either constant or varying timestep.
@@ III.1 Magnetization curve of 3d ferromagnets
The capacity of the dLLB model, as described by equations (2) and (3), to precisely forecast the experimental transition temperature $(T_{C})$ between a ferromagnetic state at low temperatures and a paramagnetic state at high temperatures is initially evaluated. For this purpose, we simulate the Curie curve of some 3d ferromagnetic simple metals as fcc-Ni, bcc-Fe, and fcc-Co, by using magnetic parameters provided in Table 1. To calculate $T_{C}$, we establish a real-space cell with a sufficient number of $N\sim 500$ interacting spins so that the influence of each neighbor on every atomic site is taken into consideration. We can monitor $N$ to account for any effect of finite size on the calculation of $T_{C}$, balancing numerical efficiency and accuracy. The equations (2) and (3) are integrated over time for each temperature $T$ within the range of 0 to 2000 Kelvin, continuing this process until the system achieves equilibrium at time $t_{eq}$. The volume average of the first moment, labelled $\mathfrak{m}$, is evaluated according to

### v2

@@ I Introduction
Tranchida et al.  [47] have formulated a rigorous and systematic statistical averaging procedure, based on functional calculus and incorporating a few microscopic material parameters. This procedure can be applied to derive set of LLB-like differential equations for the first and second moments of spin variables. The differential equation governing the dynamics of the first moment preserves the original form of the LLB equation and inherently generates a thermal longitudinal damping. Damping is calculated on the basis of variations in magnetic moments due to thermal noise, thus avoiding the need to provide prior information on longitudinal and transverse thermal damping. This equation also includes a second moment that does not factorize in a sum of first moments, suggesting that the magnetization variance may also fluctuate over time. In this model called dynamical LLB (dLLB), the damping torque that influences the precession of the magnetization, is governed by the dynamics of both the first and second moments. In return, the second moment also fluctuates in time and is itself the solution of a differential equation. The entire differential system constitutes an open hierarchy that can be closed in various ways [47] and solved numerically in a consistent manner.
@@ II Theoretical framework
To accurately simulate the ultrafast magnetization dynamics caused by laser absorption, an essential component in atomistic simulations of magnetization dynamics is needed [49]. The so-called phenomenological three-temperature model (3TM) [24] considers three interacting reservoirs, namely hot electrons, phonons and spins, and their corresponding temperatures, that exchange heat over time [50]. When a laser interacts with a metal, it generates electromagnetic waves that heat the metal. The laser energy is primarily absorbed by free electrons, leading to a rapid thermalization process at fs scale. Heavy ions do not directly absorb the laser energy, but the heated electrons transfer a small fraction of their energy to the lattice, leading to lattice heating. Then, the spins and the phonons become thermally unbalanced, and their temperatures change at a rate defined by their heat capacities and the strength of their coupling to the electron reservoir, as well as interactions among themselves.
@@
$\bm{\omega}_{i}$ $\equiv\frac{1}{\hbar}\sum_{j\neq i}J_{ij}{\bm{s}}_{j}.$ (1)
This allows the formation of a skew-symmetric matrix $\Omega$ where $\Omega_{IJ}\equiv\epsilon_{IJK}\omega_{i}^{K}$ are its components with $\epsilon_{IJK}$ being the components of the Levi-Civita antisymmetric tensor with $I,J,K\equiv x,y,z$ and $i$ being the atomic site. We also set $\Sigma_{i}\equiv\langle{\bm{S}}_{i}\otimes{\bm{S}}_{i}\rangle$ a symmetric $3\times 3$ matrix, a symmetric matrix $\Gamma_{i}\equiv{\bm{s}}_{i}\otimes\bm{s}_{i}$ and a (general) matrix $M_{i}\equiv{\bm{s}}_{i}\otimes{\bm{\omega}}_{i}$, which can be decomposed into a symmetric (resp. antisymmetric) part labelled ${M}_{i}^{S}$ (resp.${M}_{i}^{A}$). The dLLB model is reduced to the following set of equations
@@
where $\alpha$ is the dimensionless transverse Gilbert damping parameter and ${\sf{Tr}}$ stands for the trace of the matrix. In eq. (3), the commutator of two matrices is appearing as an operator that produces a matrix by contracting its internal index, thereby reducing it to a symmetric matrix, as expected. Equations (2) and (3) are numerically integrated over time by a global Runge-Kutta scheme of 4-th order, with an either constant or varying timestep. The dLLB equations are a result of statistical averaging over the noise of the well known sLLG equation. The details about the statistical averaging procedure and the model derivation are presented in our previous works [52, 53, 47].
Physically, Eq. (2) describes the dynamics of ${\bm{s}}_{i}$, the mean of the stochastic magnetic moment vector $\bm{S}_{i}$. The initial term denotes the torque associated with the precession of the magnetic moment. The subsequent term takes into account the transverse relaxation that involves a $3\times 3$ matrix ${\Sigma}_{i}$ and the final term represents the longitudinal relaxation. Equation (3) governs the dynamics of ${\Sigma}_{i}$, which represents the statistical variance of the stochastic magnetic moment $\bm{S}_{i}$. The first term describes a precession-like behavior, represented by matrix multiplication of a skew-symmetric matrix formed by the effective pulsation vector $\bm{\omega}_{i}$, and the matrix ${\Sigma}_{i}$. The second and third terms exhibit matrix transverse and longitudinal relaxation processes, respectively, similar to Equation (2), except that is for the variance.
@@ III.1 Magnetization curve of 3d ferromagnets
The capacity of the dLLB model, as described by equations (2) and (3), to precisely forecast the experimental transition temperature $(T_{C})$ between a ferromagnetic state at low temperatures and a paramagnetic state at high temperatures is initially evaluated. For this purpose, we simulate the Curie curve of some 3d ferromagnetic simple metals as fcc-Ni, bcc-Fe, and fcc-Co, by using magnetic parameters provided in Table 1. To calculate $T_{C}$, we establish a real-space cell with $N$ spins in total with periodic boundary conditions so that the influence of each neighbor on every atomic site is taken into consideration mimicking a bulk compound. The number of spins $N$ in the dLLB simulation should be sufficiently large to accurately account for all neighboring interactions. The specific size and number of atoms within each supercell are dependent on the unique characteristics of each material being investigated. For Ni and Co, we considered a supercell containing $5\times 5\times 5$ unit cells along the (x, y, z)-axes, resulting in a simulated lattice with $N=500$ atomic spins. Similarly, for Fe, we employed supercells consisting of $7\times 7\times 7$ unit cells, resulting in lattices with about $N=686$ atomic spins. The equations (2) and (3) are integrated over time for each temperature $T$ within the range of 0 to 2000 Kelvin, continuing this process until the system achieves equilibrium at time $t_{eq}$. The volume average of the first moment, labelled $\mathfrak{m}$, is evaluated according to